Grade 12Physical Chemistry

The conductivity of a saturated solution of BaSO4
is 3.06 ×10^–6 ohm–1cm–1 and its equivalent
conductance is 1.53 ohm–1cm2equiv–1. The Ksp
for BaSO4 will be

Sundarshankar Narayanan
10 Years agoGrade 12

3 Answers

Sherba
10 Years ago
If 1 mole of glucose is added to 1L of water Kb (H20) =0.512 KKg mole inverse. Boiling point of solution is what explain me with solution 
Sundarshankar Narayanan
10 Years ago
Answer given by sherba is irrelevant to the question asked here  i want full solution to this question
 
rohan
9 Years ago
Solubility product (Ksp)=(Ba2+)(SO4^2-)=C^2Molar conductivity(m)=Equivalent conductivity× equivalent factor of the electrolyte(i.e. 2 in this case) and C=Conductance/m.So,Ksp =((3.06×10^-6)/(1.53×2))^2=10^-12(answer)